# The Fundamental Parameters of Astrophysical Plasma Turbulence and its Dissipation: Nonrelativistic Limit

### v1

@@ 3 Scaling Theories for Anisotropic Plasma Turbulence
where $\alpha=0$ corresponds to the GS95 theory and $\alpha=1$ corresponds to the B06 theory. Here the turbulent fluctuations as a function of scale are characterized by the three spatial components of the wave vector relative to the local magnetic field direction: (i) the parallel wavenumber $k_{\parallel}$; and (ii) the intermediate wavenumber $k_{i}$ and (iii) the perpendicular wavenumber $k_{\perp}$, which together describe the anisotropy of the fluctuations in the plane perpendicular to the local magnetic field, $k_{i}\ll k_{\perp}$, arising from the dynamic alignment of the fluctuations (Boldyrev, 2006; Mason et al., 2006; Boldyrev et al., 2009; Perez et al., 2012) and predicting the development of current sheets at small scales. These scaling theories predict the scaling of $k_{\parallel}$, $k_{i}$, the angle of dynamic alignment $\theta$, and the amplitude of the perpendicular magnetic field fluctuations $\delta B_{\perp}$ as a function of the perpendicular wavenumber $k_{\perp}$.
@@
demonstrating that both theories predict that the turbulent fluctuations develop a scale-dependent anisotropy, whereby the fluctuations become progressively more elongated along the magnetic field as they cascade to smaller scale, eventually yielding $k_{\parallel}/k_{\perp}\ll 1$ at sufficiently small scales. This $k_{\parallel}/k_{\perp}$ anisotropy follows from the imposing of a critical balance between the linear and nonlinear timescales of the turbulence (Goldreich & Sridhar, 1995; Mallet et al., 2015). The intermediate-to-perpendicular wavevector anisotropy is given by
@@
It is worthwhile emphasizing that the idea of critical balance in the turbulent cascade does not imply that all of the turbulent energy is concentrated at the parallel wavenumber $k_{\parallel}^{cb}$ determined by critical balance. Rather, $k_{\parallel}^{cb}$ is the upper limit of parallel wavenumber, and thus generally also implies an upper limit on the linear wave frequency (since $\omega\propto k_{\parallel}$ for $\omega\ll\Omega_{i}$). Rather, the turbulent fluctuation power extends over the range $0\leq k_{\parallel}\leq k_{\parallel}^{cb}$ (Goldreich & Sridhar, 1995; Maron & Goldreich, 2001). Accounting for how the turbulent energy is distributed over this $k_{\parallel}$ range would lead to order unity or less quantitative changes in the heating rates. Since the goal of predictive turbulent heating models is generally to calculate the lowest-order, order-of-magnitude quantities for species heating rates, any effects due to the distribution of turbulent energy over $0\leq k_{\parallel}\leq k_{\parallel}^{cb}$ are likely to yield only higher-order corrections to these predictions; thus, the effect of this range of turbulent energy in $k_{\parallel}$ will be neglected in the first generation of turbulent heating models addressed here.
In closing, note that these theories for MHD turbulence are largely focused on the transport of energy in Alfvénic fluctuations, so if significant energy is transferred nonlinearly into other (non-Alfvénic ) wave modes, this may significantly alter the energy transport from the predictions shown here, with potentially qualitatively different predictions for the damping mechanisms of these non-Alfvénic turbulent fluctuations.
@@ 3.1 Exploiting Turbulent Scaling Theories to Reduce the Number of Fundamental Dimensionless Parameters
Given the large number of fundamental dimensionless parameters that characterize weakly collisional plasma turbulence, shown in Table 2, it is imperative to reduce the number of important parameters to make progress in developing predictive models of turbulent heating. Adopting idealized approximations, such as maintaining isotropic equilbrium velocity distributions for the plasma, enables one to eliminate some of the parameters in Table 2. Another approach is to exploit the MHD turbulent scaling theories described in §3 to combine several parameters into a new single parameter, an approach presented here.
@@
In Fig. 2(b) is presented the characteristic path of the anisotropic cascade of energy through $(k_{\perp},k_{\parallel})$ wavevector space for the different regimes of the turbulent cascade. There is no parallel cascade in the weak MHD turbulence regime (Sridhar & Goldreich, 1994; Galtier et al., 2000), followed by a scale-dependent anisotropy given by $k_{\parallel}\propto k_{\perp}^{2/(3+\alpha)}$ in the strong MHD turbulence regime (Goldreich & Sridhar, 1995; Boldyrev, 2006). The turbulence transitions to the strong KAW regime at $k_{\perp}\rho_{i}\sim 1$, with the anisotropy becoming stronger at $k_{\perp}\rho_{i}>1$, scaling as $k_{\parallel}\propto k_{\perp}^{1/3}$ (Cho & Lazarian, 2004; Howes et al., 2008a, 2011). As the turbulence weakens due to collisionless damping of turbulent fluctuations in the WDKT regime, the parallel cascade ceases once again (Howes et al., 2011).
@@
MHD turbulence theory predicts that the nonlinearity parameter will increase through the weak turbulent cascade until it reaches a value of unity, $\chi\sim 1$, at which point the turbulence has reached a state of critical balance in the strong turbulence regime. From that point, theory predicts that an anisotropic cascade will ensue (transferring energy to smaller scales more rapidly in the perpendicular than the parallel direction), with the parallel and perpendicular cascades balanced to maintain a nonlinearity parameter of unity, $\chi\sim 1$. Therefore, the key point in the cascade is perpendicular wavenumber at which the turbulence first reaches $\chi=1$, given by
@@
To obtain the isotropic driving wavenumber $k_{0}\rho_{i}$ from this point in wavevector space $(\left.k_{\perp}\rho_{i}\right|_{\chi=1},k_{\parallel 0})$, we need to extend the cascade backward along the critical balance line (solid black and red dashed lines) until it intersects with isotropy $k_{\parallel}=k_{\perp}$ (dotted line). From the MHD scaling theories in §3, for a critically balanced, strong turbulent cascade that is driven strongly and isotropically at $k_{0}\rho_{i}$, the anisotropy as a function of $k_{\perp}$ scales as
@@
It is worthwhile noting that the evolution of the dynamic alignment angle with perpendicular wavenumber $\theta(k_{\perp})$ is not well modeled by $k_{0}\rho_{i}$. The development of current sheets at very small scales due to dynamic alignment, and the possibility for the interruption of the turbulent cascade by magnetic reconnection (Boldyrev & Loureiro, 2017; Loureiro & Boldyrev, 2017b, a; Mallet et al., 2017b, a; Walker et al., 2018), will depend on this scaling of $\theta(k_{\perp})$. But, from a practical point of view, this issue in modeling $\theta(k_{\perp})$ while reducing ($k_{\perp 0}\rho_{i}$, $k_{\parallel 0}/k_{\perp 0}$, $\chi_{0}$) to $k_{0}\rho_{i}$ is likely to be a problem only in rare cases. First, the scaling of dynamic alignment angle $\theta\propto k_{\perp}^{-1/4}$ in B06 is rather weak, scaling as the $-1/4$ power, so only when the MHD inertial range spans many orders of magnitude in the perpendicular length scale will the development of current sheets likely have a significant impact the turbulent dynamics. Second, dynamic alignment appears to arise only in strong MHD turbulence, so it seems reasonable to take $\theta_{0}\sim 1$ radian (roughly equivalent to isotropy in the perpendicular plane) at the perpendicular scale $\left.k_{\perp}\rho_{i}\right|_{\chi=1}$ where the turbulence becomes strong. One may then estimate the evolution of that dynamic alignment angle $\theta(k_{\perp})$ from that point for the cases in which it impacts the evolution. Finally, although the development of current sheets in plasma turbulence is well established (Matthaeus & Montgomery, 1980; Meneguzzi et al., 1981; Borovsky, 2008; Uritsky et al., 2010; Osman et al., 2011; Zhdankin et al., 2012, 2013), it is not clear that the scaling theory in B06 indeed properly describes this development; an alternative explanation of the development of current sheets due to from Alfvén wave collisions exists (Howes & Nielson, 2013; Howes, 2015, 2016; Verniero et al., 2018), although a specific scaling for this alternative mechanism has not yet been proposed.
@@ 3.2 Efficiently Modeling the Small-Scale End of the Turbulent Inertial Range
The multiscale physics of astrophysical plasma turbulence presents a particular challenge for numerical modeling that covers the entire dynamic range of the turbulent cascade while simultaneously capturing the kinetic mechanisms the govern the damping of the turbulence under weakly collisional plasma conditions, in particular since physics of plasma turbulence is a inherently three-dimensional (Howes, 2015). For the development of turbulent heating models, it is critical to investigate the physical mechanisms that remove energy from the turbulence and energize the plasma particles. Fortunately, since the inertial range of turbulence is defined as the range of scales over which the turbulent energy cascades to smaller scales but the dissipation is negligible, as depicted in Fig. 1(a), it is not strictly necessary to model the the entire inertial range in order to capture the physics of the turbulent dissipation.
@@
As an example of how to set up simulations in this way, consider turbulence physically driven strongly and isotropically at an isotropic driving wavenumber of $k_{0}\rho_{i}=10^{-4}$, as depicted in Fig. 3. For a hydrogenic plasma of protons and electrons with a realistic mass ratio $m_{i}/m_{e}=1836$, the turbulent cascade extends at least down to the perpendicular scales of the electron Larmor radius $k_{\perp}\rho_{e}\sim 1$ (Sahraoui et al., 2010; Howes et al., 2011; TenBarge et al., 2013; Kiyani et al., 2015). Thus, a dynamic range of $L_{o}/\rho_{e}=k_{\perp\mbox{max}}/k_{0}=(k_{\perp\mbox{max}}\rho_{e})/(k_{0}\rho_{i})(\rho_{i}/\rho_{e})\propto(k_{0}\rho_{i})^{-1}(T_{i}/T_{e})^{1/2}(m_{i}/m_{e})^{1/2}\sim 4.3\times 10^{5}$ in each of the three spatial dimensions is necessary when $T_{i}/T_{e}\sim 1$, well beyond the capabilities of any kinetic simulation code. If the simulation code can resolve a dynamic range of $10^{3}$ in each spatial dimension, we can choose a perpendicular domain scale in the middle of the physical inertial range at $k_{\perp D}\rho_{i}=10^{-2}$ and use a reduced mass ratio $m_{i}/m_{e}=100$ so that $k_{\perp}\rho_{e}=1$ corresponds to $k_{\perp}\rho_{i}=10$ for $T_{i}/T_{e}=1$. Thus, the resolved perpendicular length scales of the simulation span $10^{-2}\leq k_{\perp}\rho_{i}\leq 10$, corresponding to $10^{-3}\leq k_{\perp}\rho_{e}\leq 1$ in terms of the electron Larmor radius $\rho_{e}$, as shown in Fig. 3.
@@ 3.3 A Reduced Parameter Space for Kinetic Plasma Turbulence
Although the physics of plasma turbulence, its dissipation, and the resulting heating of the plasma species formally depends on all ten of the general plasma and turbulence parameters $(\beta_{\parallel,i},\tau_{\parallel},A_{i},A_{e},k_{\parallel 0}\lambda_{mfp,e};k_{\perp 0}\rho_{i},k_{\parallel 0}/k_{\perp 0},\chi_{0},Z_{0}^{+}/Z_{0}^{-},E_{comp}/E_{inc})$ listed in Table 2, focusing initially on the development of completely general turbulent heating models on a ten-dimensional parameter space is unlikely to be successful. Instead, first developing predictive turbulent heating models over a reduced parameter space is likely a better strategy.
Note that kinetic temperature anisotropy instabilities generally bound the temperatures in the vicinty of isotropy, although if these instabilities are triggered, at $\beta_{i}\gg 1$, they can impact both the linear wave physics at play in the turbulent dynamics of critically balanced turbulence (Squire et al., 2016, 2017b, 2017a) and mediate the nonlocal transfer of energy directly from the large driving scales to the kinetic length scales (Arzamasskiy et al., 2023). As an initial exploration of Alfvénic turbulence and its kinetic damping mechanisms, however, we choose to focus on the simplified limit of isotropic temperatures, modifying our approach as needed for cases that violate these conditions.
@@
For many of the calculations in this study, a reduced isotropic temperature case is adopted, assuming that the equilibrium velocity distribution for each species remains well described as an isotropic Maxwellian distribution. Note that kinetic temperature anisotropy instabilities generally bound the species temperature anisotropies in the vicinity of isotropy $A_{s}\sim 1$ (Kasper et al., 2002; Hellinger et al., 2006; Štverák et al., 2008; Bale et al., 2009; Štverák et al., 2009), although if these instabilities are triggered, they can impact both the turbulent dynamics and energy transfer to particles (Squire et al., 2016, 2017b, 2017a; Arzamasskiy et al., 2023). In this isotropic temperature case, the temperature anisotropies for ions and electrons are both set to unity, $A_{i}=1$ and $A_{e}=1$. In addition, the ion plasma beta and ion-to-electron temperature ratio are redefined in terms of these isotropic temperatures, yielding the isotropic parameters $\beta_{i}=8\pi n_{i}T_{i}/B^{2}$ and $\tau=T_{i}/T_{e}$. In addition, for this reduced parameter space, we will also assume that the turbulent inertial range is sufficiently large, or the driving is sufficiently strong, that the dynamics at the perpendicular ion kinetic length scales $k_{\perp}\rho_{i}\sim 1$ are in a state of strong plasma turbulence with a nonlinearity parameter $\chi\sim 1$. Under these conditions, we can replace the three turbulent driving parameters $(k_{\perp 0}\rho_{i},k_{\parallel 0}/k_{\perp 0},\chi_{0})$ by the single isotropic driving wavenumber $k_{0}\rho_{i}$, as detailed in §3.1.
@@
We will also make two other simplfying assumptions for the isotropic temperature case: We will assume that the turbulence is balanced, with equal wave energy fluxes up and down the local magnetic field $Z_{0}^{+}/Z_{0}^{-}\sim 1$, which is a necessary condition to apply the MHD turbulence scaling relations given by (1)–(4). Additionally, based on direct spacecraft measurements of turbulence in the heliosphere which find that incompressible Alfvénic fluctuations appear to dominate turbulent dynamics in space plasmas (Belcher & Davis, 1971; Tu & Marsch, 1995; Alexandrova et al., 2008; Bruno & Carbone, 2013), as well as theoretical considerations that suggest that the dynamics of Alfvén waves dominate the turbulent cascade (Lithwick & Goldreich, 2001; Maron & Goldreich, 2001; Cho & Lazarian, 2003; Schekochihin et al., 2009; Howes, 2015), we will also assume a negligible amount of energy in compressible turbulent fluctuations, $E_{comp}/E_{inc}\ll 1$.
@@
Therefore, we propose the isotropic temperature case, a reduced parameter space for weakly collisional plasma turbulence in the limit of strong, balanced, incompressible turbulence, with a governing set of three dimensionless parameters, $(\beta_{i},\tau;k_{0}\rho_{i})$, as shown in Table 3. Although the isotropic temperature case eliminates a number of important physical effects such as temperature anisotropy instabilities, imbalanced turbulent wave energy fluxes, and compressible fluctuations, it nonetheless represents an excellent starting point for the development of refined turbulent heating models. Such models can produce theoretical predictions that can be compared to kinetic numerical simulations and spacecraft observations of weakly collisional plasma turbulence, exemplified by the phase diagram for turbulent damping mechanisms shown in Fig. 11.
@@
In terms of the fundamental parameters of the isotropic temperature case, the turbulent scaling relations (1)–(4) simplify to the following easy-to-use forms

### v2

@@ 3 Scaling Theories for Anisotropic Plasma Turbulence
where $\alpha=0$ corresponds to the GS95 theory and $\alpha=1$ corresponds to the B06 theory. Here the turbulent fluctuations as a function of scale are characterized by the three spatial components of the wave vector relative to the local magnetic field direction: (i) the parallel wavenumber $k_{\parallel}$; and (ii) the intermediate wavenumber $k_{i}$ and (iii) the perpendicular wavenumber $k_{\perp}$, which together describe the anisotropy of the fluctuations in the plane perpendicular to the local magnetic field, $k_{i}\ll k_{\perp}$, arising from the dynamic alignment of the fluctuations (Boldyrev, 2006; Mason et al., 2006; Boldyrev et al., 2009; Perez et al., 2012) and predicting the development of current sheets at small scales. These scaling theories predict the scaling of $k_{\parallel}$, $k_{i}$, the angle of dynamic alignment $\theta$ (in radians), and the amplitude of the perpendicular magnetic field fluctuations $\delta B_{\perp}$ as a function of the perpendicular wavenumber $k_{\perp}$.
@@
demonstrating that both theories predict that the turbulent fluctuations develop a scale-dependent anisotropy, whereby the fluctuations become progressively more elongated along the magnetic field as they cascade to smaller scale, eventually yielding $k_{\parallel}/k_{\perp}\ll 1$ at sufficiently small scales. This $k_{\parallel}/k_{\perp}$ anisotropy follows from the imposition of the condition of critical balance between the linear and nonlinear timescales of the turbulence (Goldreich & Sridhar, 1995; Mallet et al., 2015). The intermediate-to-perpendicular wavevector anisotropy is given by
@@
It is worthwhile emphasizing that the idea of critical balance in the turbulent cascade does not imply that all of the turbulent energy is concentrated at the parallel wavenumber $k_{\parallel}^{cb}$ determined by critical balance. Rather, $k_{\parallel}^{cb}$ is the upper limit of parallel wavenumber, and thus generally also implies an upper limit on the linear wave frequency (since $\omega\propto k_{\parallel}$ for $\omega\ll\Omega_{i}$). Rather, the turbulent fluctuation power extends over the range $0\leq k_{\parallel}\leq k_{\parallel}^{cb}$ (Goldreich & Sridhar, 1995; Maron & Goldreich, 2001). Accounting for how the turbulent energy is distributed over this $k_{\parallel}$ range would lead to order unity or less quantitative changes in the heating rates. Since the goal of predictive turbulent heating models is generally to calculate the lowest-order, order-of-magnitude quantities for species heating rates, any effects due to the distribution of turbulent energy over $0\leq k_{\parallel}\leq k_{\parallel}^{cb}$ are likely to yield only higher-order corrections to these predictions; thus, the effect of this range of turbulent energy in $k_{\parallel}$ is neglected in the first generation of turbulent heating models addressed here.
In closing, note that these theories for MHD turbulence are largely focused on the transport of energy in Alfvénic fluctuations, so if significant energy is transferred nonlinearly into other (non-Alfvénic) wave modes, this may significantly alter the energy transport from the predictions shown here, with the potential for qualitatively different predictions for the damping mechanisms of these non-Alfvénic turbulent fluctuations.
@@ 3.1 Exploiting Turbulent Scaling Theories to Reduce the Number of Fundamental Dimensionless Parameters
Given the large number of fundamental dimensionless parameters that characterize weakly collisional plasma turbulence, shown in Table 3, it is imperative to reduce the number of important parameters to make progress in developing predictive models of turbulent heating. Adopting idealized approximations, such as maintaining isotropic equilibrium velocity distributions for the plasma, enables one to eliminate some of the parameters in Table 3. Another approach is to exploit the MHD turbulent scaling theories described in §3 to combine several parameters into a new single parameter, an approach presented here.
@@
In Fig. 2(b) is presented the characteristic path of the anisotropic cascade of energy through $(k_{\perp},k_{\parallel})$ wavevector space for the different regimes of the turbulent cascade. There is no parallel cascade in the weak MHD turbulence regime (Sridhar & Goldreich, 1994; Galtier et al., 2000), followed by a scale-dependent anisotropy given by $k_{\parallel}\propto k_{\perp}^{2/(3+\alpha)}$ in the strong MHD turbulence regime (Goldreich & Sridhar, 1995; Boldyrev, 2006). The turbulence transitions to the strong KAW regime at $k_{\perp}\rho_{i}\sim 1$, with the anisotropy becoming stronger at $k_{\perp}\rho_{i}>1$, scaling as $k_{\parallel}\propto k_{\perp}^{1/3}$ (Cho & Lazarian, 2004; Howes et al., 2008a, 2011). As the turbulence weakens due to collisionless damping of turbulent fluctuations in the WDKT regime, the parallel cascade is predicted to cease once again (Howes et al., 2011).
@@
MHD turbulence theory predicts that the nonlinearity parameter will increase through the weak turbulent cascade until it reaches a value of unity, $\chi\sim 1$, at which point the turbulence has reached a state of critical balance in the strong turbulence regime. From that point, theory predicts that an anisotropic cascade will ensue (transferring energy to smaller scales more rapidly in the perpendicular than the parallel direction), with the parallel and perpendicular cascades balanced to maintain a nonlinearity parameter of unity, $\chi\sim 1$. Therefore, the key point in the cascade is the perpendicular wavenumber at which the turbulence first reaches $\chi=1$, given by
@@
To obtain the isotropic driving wavenumber $k_{0}\rho_{i}$ from this point in wavevector space $(\left.k_{\perp}\rho_{i}\right|_{\chi=1},k_{\parallel 0})$, we need to extend the cascade backward along the critical balance line (red dashed lines) until it intersects with isotropy $k_{\parallel}=k_{\perp}$ (dotted line). From the MHD scaling theories in §3, for a critically balanced, strong turbulent cascade that is driven strongly and isotropically at $k_{0}\rho_{i}$, the anisotropy as a function of $k_{\perp}$ scales as
@@
It is worthwhile noting that the evolution of the dynamic alignment angle with perpendicular wavenumber $\theta(k_{\perp})$ is not well modeled by $k_{0}\rho_{i}$. The development of current sheets at very small scales due to dynamic alignment, and the possibility for the interruption of the turbulent cascade by magnetic reconnection (Boldyrev & Loureiro, 2017; Loureiro & Boldyrev, 2017b, a; Mallet et al., 2017b, a; Walker et al., 2018), will depend on this scaling of $\theta(k_{\perp})$. But, from a practical point of view, this issue in modeling $\theta(k_{\perp})$ while reducing ($k_{\perp 0}\rho_{i}$, $k_{\parallel 0}/k_{\perp 0}$, $\chi_{0}$) to $k_{0}\rho_{i}$ is likely to be a problem only in rare cases, for the following reasons. First, the scaling of dynamic alignment angle $\theta\propto k_{\perp}^{-1/4}$ in B06 is rather weak, scaling as the $-1/4$ power, so only when the MHD inertial range spans many orders of magnitude in the perpendicular length scale will the development of current sheets likely have a significant impact on the turbulent dynamics. Second, dynamic alignment appears to arise only in strong MHD turbulence, so it seems reasonable to take $\theta_{0}\sim 1$ radian (roughly equivalent to isotropy in the perpendicular plane) at the perpendicular scale $\left.k_{\perp}\rho_{i}\right|_{\chi=1}$ where the turbulence becomes strong. One may then estimate the evolution of that dynamic alignment angle $\theta(k_{\perp})$ from that point for the cases in which it impacts the evolution. Finally, although the development of current sheets in plasma turbulence is well established (Matthaeus & Montgomery, 1980; Meneguzzi et al., 1981; Borovsky, 2008; Uritsky et al., 2010; Osman et al., 2011; Zhdankin et al., 2012, 2013), it is not clear that the scaling theory in B06 indeed properly describes this development; an alternative explanation for the development of current sheets arising self-consistently from Alfvén wave collisions exists (Howes, 2015, 2016; Verniero et al., 2018), although a specific scaling for this alternative mechanism has not yet been proposed.
@@ 3.2 Efficiently Modeling the Small-Scale End of the Turbulent Inertial Range
The multiscale physics of astrophysical plasma turbulence presents a particular challenge for numerical modeling that covers the entire dynamic range of the turbulent cascade while simultaneously capturing the kinetic mechanisms that govern the damping of the turbulence under weakly collisional plasma conditions, in particular since the physics of plasma turbulence is inherently three-dimensional (Howes, 2015). For the development of turbulent heating models, it is critical to investigate the physical mechanisms that remove energy from the turbulence and energize the plasma particles. Fortunately, since the inertial range of turbulence is defined as the range of scales over which the turbulent energy cascades to smaller scales but the dissipation is negligible, as depicted in Fig. 1(a), it is not strictly necessary to model the the entire inertial range in order to capture the physics of the turbulent dissipation.
@@
As an example of how to set up simulations in this way, consider turbulence physically driven strongly and isotropically at an isotropic driving wavenumber of $k_{0}\rho_{i}=10^{-4}$, as depicted in Fig. 3. For a hydrogenic plasma of protons and electrons with a realistic mass ratio $m_{i}/m_{e}=1836$, the turbulent cascade extends at least down to the perpendicular scales of the electron Larmor radius $k_{\perp}\rho_{e}\sim 1$ (Sahraoui et al., 2010; Howes et al., 2011; TenBarge et al., 2013; Kiyani et al., 2015). Thus, a dynamic range of $L_{0}/\rho_{e}=k_{\perp\mbox{max}}/k_{0}=[(k_{\perp\mbox{max}}\rho_{e})/(k_{0}\rho_{i})](\rho_{i}/\rho_{e})\propto(k_{0}\rho_{i})^{-1}(T_{i}/T_{e})^{1/2}(m_{i}/m_{e})^{1/2}\sim 4.3\times 10^{5}$ in each of the three spatial dimensions is necessary when $T_{i}/T_{e}\sim 1$, well beyond the capabilities of any kinetic simulation code. If the simulation code can resolve a dynamic range of $10^{3}$ in each spatial dimension, we can choose a perpendicular domain scale in the middle of the physical inertial range at $k_{\perp D}\rho_{i}=10^{-2}$ and use a reduced mass ratio $m_{i}/m_{e}=100$ so that $k_{\perp}\rho_{e}=1$ corresponds to $k_{\perp}\rho_{i}=10$ for $T_{i}/T_{e}=1$. Thus, the resolved perpendicular length scales of the simulation span $10^{-2}\leq k_{\perp}\rho_{i}\leq 10$, corresponding to $10^{-3}\leq k_{\perp}\rho_{e}\leq 1$ in terms of the electron Larmor radius $\rho_{e}$, as shown in Fig. 3.
@@ 3.3 A Reduced Parameter Space for Kinetic Plasma Turbulence
Although the physics of plasma turbulence, its dissipation, and the resulting heating of the plasma species formally depends on all ten of the general plasma and turbulence parameters $(\beta_{\parallel,i},\tau_{\parallel},A_{i},A_{e},k_{\parallel 0}\lambda_{mfp,e};k_{\perp 0}\rho_{i},k_{\parallel 0}/k_{\perp 0},\chi_{0},Z_{0}^{+}/Z_{0}^{-},E_{comp}/E_{inc})$ listed in Table 3, focusing initially on the development of completely general turbulent heating models on a ten-dimensional parameter space is unlikely to be successful. Instead, first developing predictive turbulent heating models over a reduced parameter space is almost certainly a better strategy. As an initial exploration of Alfvénic turbulence and its kinetic damping mechanisms, we choose to focus on the simplified limit of isotropic temperatures, modifying our approach as needed for cases that violate these conditions.
@@
For many of the calculations in this study, a reduced isotropic temperature case is adopted, assuming that the equilibrium velocity distribution for each species remains well described as an isotropic Maxwellian distribution. Note that kinetic temperature anisotropy instabilities generally bound the species temperature anisotropies in the vicinity of isotropy $A_{s}\sim 1$ (Kasper et al., 2002; Hellinger et al., 2006; Štverák et al., 2008; Bale et al., 2009; Štverák et al., 2009), although if these instabilities are triggered, particularly at $\beta_{i}\gg 1$, they can impact both the linear wave physics at play in the turbulent dynamics of critically balanced turbulence (Squire et al., 2016, 2017a, 2017b) and mediate the nonlocal transfer of energy directly from the large driving scales to the kinetic length scales (Arzamasskiy et al., 2023).
To define a reduced parameter space, we first restrict the temperature anisotropies for ions and electrons to unity, $A_{i}=1$ and $A_{e}=1$, so that $T_{\perp,i}=T_{\parallel,i}\equiv T_{i}$ and $T_{\perp,e}=T_{\parallel,e}\equiv T_{e}$. With this idealization, the ion plasma beta and ion-to-electron temperature ratio are redefined in terms of these isotropic temperatures $T_{i}$ and $T_{e}$, yielding the isotropic parameters $\beta_{i}=8\pi n_{i}T_{i}/B^{2}$ and $\tau=T_{i}/T_{e}$. In addition, for this reduced parameter space, we will also assume that the turbulent inertial range is sufficiently large, or the driving is sufficiently strong, that the dynamics at the perpendicular ion kinetic length scales $k_{\perp}\rho_{i}\sim 1$ are in a state of strong plasma turbulence with a nonlinearity parameter $\chi\sim 1$. Under these conditions, we can replace the three turbulent driving parameters $(k_{\perp 0}\rho_{i},k_{\parallel 0}/k_{\perp 0},\chi_{0})$ by the single isotropic driving wavenumber $k_{0}\rho_{i}$, as detailed in §3.1.
@@
We will also make two other simplifying assumptions for this reduced, isotropic temperature case. We will assume that the turbulence is balanced, with approximately equal wave energy fluxes up and down the local magnetic field $Z_{0}^{+}/Z_{0}^{-}\sim 1$, which is a necessary condition to apply the MHD turbulence scaling relations given by (1)–(4). Additionally, based on direct spacecraft measurements of turbulence in the heliosphere which find that incompressible Alfvénic fluctuations appear to dominate turbulent dynamics in space plasmas (Belcher & Davis, 1971; Tu & Marsch, 1995; Alexandrova et al., 2008; Bruno & Carbone, 2013), as well as theoretical considerations that suggest that the dynamics of Alfvén waves dominate the turbulent cascade (Lithwick & Goldreich, 2001; Maron & Goldreich, 2001; Cho & Lazarian, 2003; Schekochihin et al., 2009; Howes, 2015), we will also assume a negligible amount of energy in compressible turbulent fluctuations at the driving scale, $E_{comp}/E_{inc}\ll 1$.
@@
Therefore, we propose the isotropic temperature case, a reduced parameter space for weakly collisional plasma turbulence in the limit of strong, balanced, Alfvénic turbulence, with a governing set of three dimensionless parameters, $(\beta_{i},\tau;k_{0}\rho_{i})$, as shown in Table 4. Although the isotropic temperature case eliminates a number of important physical effects such as temperature anisotropy instabilities, imbalanced turbulent wave energy fluxes, and compressible fluctuations, it nonetheless represents an excellent starting point for the development of refined turbulent heating models. Such models can produce theoretical predictions that can be compared to kinetic numerical simulations and spacecraft observations of weakly collisional plasma turbulence, exemplified by the phase diagram for turbulent damping mechanisms shown in Fig. 12.
@@
In terms of the fundamental parameters of the isotropic temperature case, the MHD turbulent scaling relations (1)–(4) simplify to the following easy-to-use forms